hartnn (hartnn):

I would like to ask a similar(not same) question what Directrix has asked. HOW SHOULD BE OUR RESPONSE ? And do we report Response 2?

hartnn (hartnn):

let me also give an example : Question : 4(x+9) = 2(x+20), solve for x. Response 1: x= 2 <-----direct answer, obviously not allowed here... ------------------------------------------------------ Response 2: 4x+36=2x+40 4x-2x=40-36 2x=4 x=2 Now this is complete solution showing all steps. Can we do this here ? Would it be violation of code of conduct? ------------------------------------------------------- Response 3: Can you distribute using a(b+c)=ab+ac ? asker:Yes. So, what will be 4(x+9) ? .... and so on. Now this is the way , i think, the asker should be guided, as this enables learning process.

hartnn (hartnn):

I had following situation more than once. I was explaining the asker the steps to solve the question one by one, and the asker was also responding correctly. I had explained only half the solution(or few steps), and suddenly some other OS user posts the entire solution with all steps and final answer. The asker closed the question and went to ask new question. What should i have done here? Report that OS user?(or i cannot because he gave all steps and not just final answer) or should i have given entire solution at once in the beginning itself ??

OpenStudy (anonymous):

That happens frequently enough to be annoying....

Directrix (directrix):

@hartnn I viewed this OS video on the OS Facebook site. The helper owl gave what could be thought of as the "entire solution" to the asker owl without calling upon the asker owl for input. Interesting. https://www.facebook.com/#!/photo.php?v=546254509750&set=vb.85705490637&type=2&theater

OpenStudy (anonymous):

I usually request feedback because quite often the op disappears half way through (was it something i said..?)

mathslover (mathslover):

That is known as "answer snipping" . But, if the user annoys with any bad behavior or ... by typing the entire solution or... giving direct answer then it is wrong. If the user corrects you or gives suggestion in any respect , then it cant' be referred as wrong unless he/she annoys you. This "answer snipping" was very more earlier but now it is being improved. @hartnn basically, as per my views there are two ways: i) To instruct the user kindly to not interrupt. ii) "Report abuse" is the ultimate option for us.

hartnn (hartnn):

after posting the entire solution, that user goes immediately to other question for posting the solution...and never returns....same observation,everytime.

mathslover (mathslover):

hmn, in that case Report Abuse is the best option.

OpenStudy (anonymous):

1) The mods should really think twice before warning/banning - read whether the asker was so SO, SO helpless, and SO hapless that the Answerer just had to finish by giving it all. In this case THERE IS NO OTHER WAY !!! 2) If they really want coherence then they MUST STATE EXPLICITLY TO ALL: "No complete solutions are provided here" " Please medal answerers for central piece of solutions."

mathslover (mathslover):

or in the worst case , message mods (not admins) regarding the behavior of the user.

mathslover (mathslover):

But still, I wish admins to appoint some mods at this time standard. NO MODS present at present.

OpenStudy (anonymous):

And as we actually know , all of us know that from experince \[\Large \text{Absolute Power Corrupts absolutely}\] the criticism of mods should be A) In interactive way - giving chance to defend yourself B) There should be a rare Higher authority TO APPEAL

OpenStudy (anonymous):

Guys - you are double lucky 9*9 AND green. Why are you so quiet?... speak please

mathslover (mathslover):

@Mikael who said that mods have absolute power?

OpenStudy (anonymous):

Where is the appeal button ? \smirk \sarc

mathslover (mathslover):

In my opinion, Admins should think of distributing powers among good users / responsible ones. And, mods don't have supreme power, ADMINS are the supreme.

OpenStudy (anonymous):

This is called OPEN study, so where is the open FORUM of discussion

hartnn (hartnn):

i like the idea of distribution of powers...

OpenStudy (anonymous):

Why not let the greens be the senate ?

mathslover (mathslover):

@Mikael that's not possible , since some are users which are expert but don't get enough time to helps others, and some are there who dont' know much but are green. Hence we can not be "senate".

OpenStudy (anonymous):

Senate can be some sub-set of greens. But they definitely should not include the executive - not the purples.

mathslover (mathslover):

This is an important issue for OpenStudy.

OpenStudy (anonymous):

Well idea has been "incepted" the senate of "Green-bloods" - a Magna Carta of kinds. The double chamber structure will come later...

OpenStudy (anonymous):

Aand a Dual- complement of the above: Is any editorial attention being paid to simply answer-gatherers ? People who ask, but who actively oppose explanations, only FULL solutions. I'm observing such behaviour frequently

Directrix (directrix):

When we try to help someone and that person will not respond (or goes offline), then is it better to leave the question hanging incomplete or is it wiser to complete the solution? With a complete solution, somebody somewhere might benefit from having a sample solution as a prototype to follow in working problems of the same ilk. With a hanging solution, the person might return and respond later. These problems and solutions (the entire thread, really) leaves OpenStudy and goes out into WebWorld and can be seen by anyone. Theoretically so, I suppose.

OpenStudy (anonymous):

@Directr?ix @hartnn @mathslover What do YOU do when the user wants it "on a silver platter" OR NOTHING LESS

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

@mathslover Wow, you are correct. There are no mods here now.

Directrix (directrix):

@Mikael I flat out tell the Asker that the site does not allow solutions to be plastered on the thread without some input from the Asker, adding that we are to engage the Asker in the solution process. Because some Answerers give out answers, the Asker many times ignores me and waits for an answer "gift" from someone else. Other times, the Asker goes offline and accesses the site as a "ghost observer" and waits to see what happens. Today, I spent an hour with two people working two Algebra One problems. Most of that time was spent in begging for input from the Asker. Like the person who said he had received one warning but didn't know for which problem he didn't show work, so have I received one warning. No idea when I did not show work or somehow failed to meet the Code of "Guidance." When an Asker does not give an immediate reply, sometimes the Asker is timid, here for the first time, scared, or feeling inadequate or other things. So, I try to wait until the Asker feels comfortable in posting. Many times, the Asker does not post anything. I get that nobody wants to look bad but this is a helping site. As I tell the Asker, you can delete a post if you don't like what you wrote. This is why guidance on "guiding to a solution" guidelines would be wonderful. I want to carry out the Administrators' specific mission for OS if I could understand what it is. But, they will have to write with the precision normally found in mathematics.

hartnn (hartnn):

ditto what he said ^^

mathslover (mathslover):

It will be also better if you invite senior members like @sasogeek ...

OpenStudy (sasogeek):

how should a response be? 1. A one reply to a question as an answer is not enough, and is potentially dangerous 2. AT LEAST first find out what the asker already knows about the solution, what they've already tried, if they've done any research..... start a conversation 3. make sure the asker UNDERSTANDS HOW TO find the answer. however you do that is up to you, but engage the asker. 4. IF the asker permits, and you can, test the asker's new found knowledge you imparted. if the asker wants just the answer or is not ready to be engaged in the solution process, ignore them. you might want to read my post in Directrix's thread :)

mathslover (mathslover):

@sasogeek agreed but what in the case if "answer snipping" is done? Reporting? But that report comes to the notification to the mods/admins late and the user (who is doing answer snipping) continues to carry his process.

hartnn (hartnn):

Thanks for the reply sasogeek. I appreciate it. So, when i am in the middle of an explanation and someone posts the entire solution(answer snipping) , what I will do is REPORT that person (i already report many people everyday! increasing the workload of moderators....) and if the asker intends then i will continue with an explanation else just go to next problem. I would still like opinion on this by other mods/admins/ambassadors/os users. So i will keep this open for a while.

OpenStudy (sasogeek):

certainly moderating the site and taking actions towards reports is not an easy task, but what you can also certainly do to help is make the report first, if you can warn and rebuke the answer sniper in the same thread, and continue your help with the asker. you can't control everything to be what you want it to be in a perfect society so sometimes the asker may leave after getting an answer snipe, but when they leave, they harm their own selves, the answer sniper gets banned, and life moves on :) if the asker wants understanding, they will stay and learn from your guidance.

OpenStudy (shadowfiend):

Sasogeek pretty much covers it. Please do report people who interrupt you and give the answer, work shown or not. “B) There should be a rare Higher authority TO APPEAL” Every warning includes a note that you should email abuse AT openstudy.com if you have any questions.

mathslover (mathslover):

But @shadowfiend there is a user who is answering questions directly ... I had reported him many times (regarding different complaints) . I don't think that any action is taken for him.

mathslover (mathslover):

I am not blaming your decisions but like just confirming : i) whether you received my report ii) or you didn't get any wrong there.

OpenStudy (shadowfiend):

Actions have likely been taken behind the scenes. Do email abuse AT openstudy.com with whatever issues you have and we'll look into it though. I personally don't get a chance to read every report, that's why we have a team of moderators :)

mathslover (mathslover):

Is email abuse and report abuse the same thing?

OpenStudy (shadowfiend):

No. You shouldn't email abuse AT openstudy.com unless you feel you're being repeatedly ignored, or you have a question about a warning or suspension. Report Abuse is the proper path 99.99% of the time.

mathslover (mathslover):

OK but how do I confirm myself that I am "repeatedly ignored" ? I even don't know exactly whether the user has faced any action by mods or not. Should I talk to an admin in that case?

hero (hero):

@hartnn I'm not a fan of giving answers either, however, some users need to actually see the process of how the problem is solved as well. I'm not a fan of giving direct answers, however, I realize that some users do need to see the process, and yes, actually some users may even need to know what the correct answer is. This way, they know for certain what it is without second guessing themselves. I am not a fan of users who come online and post test questions and if that is the case, I do report users. I also report users who give a direct answer without explanation. For many of the computational questions, I usually do something similar to Response 2, but for the most part, I do explain the steps. Some steps are obvious, but I think a good compromise for a response 2 is to maybe do something like this: Response 2: 4x+36=2x+40 4x-2x=40-36 2x=4 x=? Show the process and leave a question mark next to the end result is probably something that can be acceptable for the computational questions. I support showing the students the process. I know some students on here are looking for just answers, but I will not allow that to stop me from helping students who actually want to learn. That's all I have to say about it.

hartnn (hartnn):

but i would rather not give them many steps at a time.....users may require all steps sometimes, but they need to mention that. like when we do 4x-2x=40-36 2x=4 this they can't do on their own....so we need to make them understand what and why is that needed to be done...if they don't do it by themselves, they will not be able to do it for next problems...

hartnn (hartnn):

sorry, i meant this : 4x+36=2x+40 4x-2x=40-36

hero (hero):

You mean, some users may not be able to do that on their own. That's the purpose of explaining steps of course. If I see a user posting multiple questions, I definitely expect for them to be learning the process, and when they post similar questions after they have shown the process, I expect them to demonstrate their learning. But if a student continues to post questions without regard to that process, I suddenly become dis-interested in continuing to help that student. I think each student should be handled on a case by case basis. I don't think its fair to say that every student doesn't get the solving process because that just simply isn't true. But I understand where you're coming from. Also understand that we all have our different approaches to helping students, so when you have an opportunity to use your particular approach, by all means, use it. I'm sure the student as well as other users will appreciate it.

hartnn (hartnn):

i so agree on this : " I think each student should be handled on a case by case basis."